The perimeter of a rectangle is 110 feet. Find the length and width if the width is 5 less than the length.
Answer by waynest(281)

width = L -5
length = L
perimeter = 110
2L + 2(L - 5) = 110
2L + 2L - 10 = 110
4L - 10 + 10 = 110 + 10
4L = 120
4L/4 = 120/4
L = 30
check:
2(30) + 2(30 -5) = 110
60 + 2(25) = 110
60 + 50 = 110
110 = 110